What volume did a helium-filled balloon have at 26.8 °C and 1.87 atm if its new volume was 63.2 mL at 2.22 atm and 9.6°C? O a. 79.6 O b. 209 O c. 70.7 O d. 56.5


1
Expert's answer
2022-05-26T18:49:03-0400

In an isolated system, the mole number remains the same:

"n_1=n_2"


In the ideal gas law, the mole number is related to the temperature T, volume V, pressure P and gas constant R:

"pV=nRT"

"n=\frac{pV}{RT}"


Thus:

"\frac{p_1V_1}{RT_1} = \frac{p_2V_2}{RT_2}"


The initial volume V1 can be deduced from the last equation:

"V_1 = \frac{p_2V_2T_1}{T_2p_1} = \frac{2.22 \times 63.2 \times 26.8}{9.6 \times 1.87} = 209 ml"

Answer: b 209 ml